As an analyst at Walmart, your job is to observe consumer transactions and help make informed

economic decisions. On Day 1 when the per-unit price of bread is $5, 25 are sold. On Day 2 when the

per-unit price of bread is $7, 15 are sold. Calculate the elasticity of demand.

Please show the work to how you got the answer

(a) -2/9

(b) -1

(c) -3/2

(d) -9/4

(e) -5/7
Answer #1

ANSWER:

Elasticity of demand = % change in quantity demanded / % change in price

% change in quantity demanded = (q2 - q1) / (q2 + q1) / 2 = (15 - 25) / (15 + 25) / 2 = -10 / 20 = -0.5

% chang ein price = (p2 - p1) / (p2 + p1) / 2 = (7 - 5) / (7 + 5) / 2 = 2 / 6 = 1 / 3 = 0.33

Elasticity of demand = - 0.5 / 0.33 = -3 / 2

so the correct answer is option c.